OpenStudy (anonymous):

a right angled triangle ABC, right angled at B , prove that sinA / cosA = tan A and cosA / sinA = cot A

OpenStudy (anonymous):

 use the definitions

OpenStudy (anonymous):

tan A = (a/c) = (a/b) / (c/b)

OpenStudy (anonymous):

LHS= sinA=BC/AC cos A = AB/AC sinA/cosA=(BC/A)/(AB/AC)=BC/AB RHS= tanA=BC/AB hence, LHS= RHS

OpenStudy (anonymous):

hmm...yes i have done sme method as @susan3 wrote